A tugboat goes 224 miles upstream in 16 hours. The return trip downstream takes 8 hours. Find the speed of the tugboat without a current and the speed of the current.
The speed of the tugboat without a current is__________miles/hour. (If needed, round to 2 decimal places)
The speed of the current is________miles/hour. (If needed, round to 2 decimal places)



Answer by Boreal(15235)

speed=s
current=c
speed (total) upstream is 224 miles/16 hours or 14 mph. The speed downstream is 224/8=28 mph
s-c=14 (units in mph)
s+c=28
2s=42
s=21 mph in calm water. This is the speed of the boat
c=7 mph. This is the speed of the current.